### v1

@@ 3.3 Main Result
As we discuss in the next section, the above result is an extension of earlier results of [4] and the references therein. We additionally show that when the MFC problem has the widely employed separable structure, the resulting MFG is a standard one which depends only on the distribution of the state process and not the control.
4 Discussion and Special Cases
4.1 Separable Running Cost
We first investigate separable cost functions. Assume that for suitable functions,
$F_{0}:[0,T]\times{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},\ \ F_{1}:[0,T]\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},\ \ G_{0}:{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},$
where $\mu=\nu\circ\Pi^{-1}$ and, as before, $\Pi:{\mathscr{D}}_{n}\times U\mapsto{\mathscr{D}}_{n}$ is the projection on the first variable. This setting corresponds to a standard mean-field game, and in most applications $F_{0}=\tfrac{1}{2}|a|^{2}$. Then, the costs $f(t,x,\mu,a)$, $g(t,\nu)$ given in (1.4) depend only on $\mu=\nu\circ\Pi^{-1}$ and we have
$f(t,x,\nu,a)=F_{0}(t,x,a)+\delta_{\mu}F_{1}(t,\mu)(x)+F_{1}(t,\mu),\quad g(x,\mu)=\delta_{\mu}G(\mu)(x)+G(\mu).$
For a given flow $t\mapsto\mu_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n})$, the terms $F_{1}(t,\mu_{t})$ and $G(\mu_{T})$ are independent of the control. Therefore, they do not influence the control problems and without loss of generality, we omit them in the definitions of the potential MFG.
The optimization of the representative agent in the MFG is to minimize
$J_{g}(X_{0},\alpha,\mu)=\mathbb{E}\left[\int_{0}^{T}\left[F_{0}(t,X^{\alpha}_{\cdot\land t},\alpha_{t})+\delta_{\mu}F_{1}(t,\mu_{t})(X^{\alpha}_{\cdot\land t})\right]\,\mathrm{d}t+\delta_{\mu}G(\mu_{T})(X^{\alpha}_{T})\right],$ (4.2)
The following is a direct consequence of Theorem 3.6 and Theorem 7.2.
Corollary 4.1.
Suppose Assumptions 2.3, 3.5 and 7.1 hold, and assume that $\alpha^{*}\in{\mathscr{A}}$ is a minimizer of the MFC problem (3.1) without common noise and with running cost given by (4.1). Then, with $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t})$ the pair $(\alpha^{*},\mu^{*})$ is a Nash equilibrium for the potential MFG with the cost functional (4.2).
4.2 Markovian Cost Without Jumps
In addition to separability of the cost functionals, we now assume that $\lambda=0$. Then, the state process $X^{\alpha}$ is continuous in time, and the invariance is proved in [26, Proposition 2.4]. Following [4], we also assume that the cost functions in (4.1) are Markovian, i.e. at time $t\in[0,T]$ the dependence on the state $X$ is only through the current position $X_{t}$. Hence, the payoff functional (3.1) has the form,
$J(\alpha,X_{0})=\mathbb{E}\left[\int_{0}^{T}\left(\tilde{F}_{0}(t,X^{\alpha}_{t},\alpha_{t})+\tilde{F}_{1}(t,{\mathcal{L}}(X^{\alpha}_{t}))\right)\,\mathrm{d}t\right]+\tilde{G}({\mathcal{L}}(X^{\alpha}_{T})),$ (4.3)
for some functions $\tilde{F}_{0}$, $\tilde{F}_{1}$ and $\tilde{G}$.
The corresponding MFG cost functional is given by,
$J_{g}(X_{0},\alpha,\mu)=\mathbb{E}\left[\int_{0}^{T}\left[\tilde{F}_{0}(t,X^{\alpha}_{t},\alpha_{t})+\delta_{\mu}\tilde{F}_{1}(t,\mu_{t})(X^{\alpha}_{t})\right]\,\mathrm{d}t+\delta_{\mu}\tilde{G}(\mu_{T})(X^{\alpha}_{T})\right],$
This is exactly the MFG studied in many of the earlier studies such as [4]. In particular, it is a standard MFG as the mean-field interaction is only through the law of the current state and not the control. Then, the measure flow takes values in ${\mathscr{P}}_{2}(\mathbb{R}^{n})$ instead of ${\mathscr{P}}_{2}({\mathscr{D}}_{n})$ or ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$. Further, by the continuity of $t\mapsto X^{\alpha}_{t}$, the flow $t\mapsto\mu_{t}\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ is continuous, and the Corollary 4.1 holds with a continuous deterministic flow $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{t})$.
4.3 Deterministic and Stochastic Formulations
Although the formulation of [4] is seemingly different than that of ours, they are equivalent and their results are included in our framework. Indeed, control problems on the space of measures have been studied in varying settings. In particular, [25, 26] provide a comprehensive overview of different stochastic formulations of the problem. In the absence of common noise and with Markovian dynamics, these problems also admit deterministic formulations that directly control the evolution of laws via the Fokker-Planck equation [3, 4, 5, 16, 49, 51, 57, 58].
$\mathrm{d}X^{\alpha}_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\mathrm{d}W_{t}.$
In this setting these two formulations are the followings.
  1. (1)
  2. (2)
    $\int_{0}^{T}[\,\int_{\mathbb{R}^{n}}\tilde{F}_{0}(t,x,\alpha(t,x))\,\mu^{\alpha}_{t}(\mathrm{d}x)\ +\ \tilde{F}_{1}(t,\mu^{\alpha}_{t})]\,\mathrm{d}t+\tilde{G}(\mu^{\alpha}_{T}),$
    over all measurable feedback functions $\alpha:[0,T]\times\mathbb{R}^{n}\mapsto U$, where the state $\mu^{\alpha}\in{\mathscr{P}}(\mathbb{R}^{d})$ solves the Fokker-Planck equation:
The equivalence between between these two formulations is proved using the superposition principle that relates a solution of a Fokker–Planck equation to a weak solution of a SDE, see for example [45] which establish this fact under much more general dynamics. Starting from the deterministic formulation, this allows us to define a weak control for the stochastic problem in feedback form with the probability space being part of the control. One then appeals to a mimicking theorem to establish the equivalence of weak control processes and weak controls in feedback form. Additional arguments such as boundedness of feedback controls are utilized to obtain a strong control on a fixed probability space. In the present setting, a detailed implementation of these steps are given in [16, Proposition 2.2].
@@ 4.3 Deterministic and Stochastic Formulations
The main idea is to compose the optimal control $\alpha^{*}$ with an arbitrary control $\alpha$ through a biased independent coin-flip. To achieve this, we extend the original probabilistic structure by adding an initial Bernoulli random variable, and then we use either one of the controls depending on its outcome. This provides linearity in the distribution variables and we obtain the result by essentially differentiating with respect to success probability of the coin-flip. The assumed invariance principle ensures that $\alpha^{*}$ is still a minimizer of the extended problem.
@@ 5.1 Preliminaries
Then, for any $Z=Z(\omega,x)$, $\hat{Z}=Z\otimes Z$ and $Z=Z^{(0)}\otimes Z^{(1)}$ where $Z^{(x)}(\omega)=Z(\omega,x)$.
@@
This defines a family of probability bases $\gamma^{\delta}=(\hat{\Omega},\hat{\mathscr{F}},\mathbb{P}^{\delta},\hat{\mathbb{F}},\hat{\mathscr{W}},\hat{\mathscr{N}})$ indexed by $\delta\in(0,1)$. As $\mathbb{P}^{\delta}$ is equivalent to $\hat{\mathbb{P}}$, we have $\hat{\mathscr{A}}:={\mathscr{A}}(\hat{\gamma})={\mathscr{A}}(\gamma^{\delta})$. For $\beta\in\hat{\mathscr{A}}$, the projected controls $\beta^{(0)},\beta^{(1)}$ are in ${\mathscr{A}}$ and by definition $\beta_{t}=\beta^{(0)}_{t}\otimes\beta^{(1)}_{t}$. Let $Y^{\beta}$ be the solution to $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$, and $X^{(0)}:=X^{\beta^{(0)}}$, $X^{(1)}:=X^{\beta^{(1)}}$ be the controlled state processes solving the equation SDE$(X_{0},\alpha)$ with controls $\beta^{(0)}$ and $\beta^{(1)}$, respectively. Further, we set,
@@
where the final equality follows by properties of the stochastic integral, see for example [41, Exercise 2.30,Chapter 3.2]. Hence, $I_{t}(\beta)=I_{t}(\alpha^{(0)})\otimes I_{t}(\alpha^{(1)})$ a.s. A similar computation of all other four integrals in $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$ implies that $X^{(0)}\otimes X^{(1)}$ is a solution of $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$. Then by strong uniqueness, we conclude that $Y^{\beta}=X^{(0)}\otimes X^{(1)}$.
Let $\xi:(\omega,x)\in\hat{\Omega}\mapsto\xi(\omega,x)=x$ be the canonical projection. For $t\in[0,T]$ and a Borel set $S\subset{\mathscr{D}}_{n}\times U$, the followings hold a.s.,
$\nu^{\beta,\delta}(S)$ $=\mathbb{P}^{\beta}((Y^{\beta}_{\cdot\land t},\beta_{t})\in S\,|\,\hat{\mathscr{G}})$
@@
$\nu^{\beta,\delta}(S)(\omega,x)$ $=\mathbb{P}^{\delta}(\xi=0)\ \mathbb{P}((X^{(0)}_{\cdot\land t},\alpha^{(0)}_{t})\in S\,|\,{\mathscr{G}})+\mathbb{P}^{\delta}(\xi=1)\ \mathbb{P}((X^{(1)}_{\cdot\land t},\alpha^{(1)}_{t})\in S\,|\,{\mathscr{G}})$
$=\delta\,\nu^{(0)}(S)(\omega)+(1-\delta)\,\nu^{(1)}(S)(\omega).$
@@ 5.2 Completion of the Proof
$0\leq\lim_{\delta\downarrow 0}\frac{1}{\delta}\left(J_{\delta}(\hat{X}_{0},\beta)-J(X_{0},\alpha^{*})\right)=J_{g}(\alpha,X_{0},\nu^{*})-J_{g}(X_{0},\alpha^{*},\nu^{*}).$
In particular, $\alpha^{*}$ minimizes $J_{g}(\alpha,X_{0},\nu^{*})$ over all $\alpha\in{\mathscr{A}}$.
@@
$\hskip 50.0pt=\mathbb{E}\left[\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\land t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\,|\,{\mathscr{G}}\,\right]\right]$
@@
We have shown that $\alpha^{*}$ minimizes $J_{g}(X_{0},\alpha,\nu^{*})$ and by its definition $\nu^{*}_{t}={\mathcal{L}}((X^{*}_{\wedge t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$. Hence, $(\alpha^{*},\nu^{*})$ is a Nash equilibrium of MFG of controls.
@@
We use the invariance property only in (5.3). Therefore it suffices to require that value functions on $\gamma$ and its extension $\hat{\gamma}$ have the same value functions. In fact, if the probability basis $\gamma$ is rich enough in the sense defined in [20, Section 2.1.3] and $X_{0}$ is deterministic, one can directly prove that the extension to $\hat{\gamma}$ does not change the value.

### v2

@@ 3.3 Main Result
As we discuss in Section 5.4, our above result is an extension of the connection proposed in [6] and the references therein. We additionally show that when the MFC problem has the widely employed separable structure, the resulting MFG is a standard one which depends only on the distribution of the state process and not the control.
3.4 Law-Dependent Dynamics
This section discusses a variation of our approach to MFC problems in which the controlled state dynamics are allowed to depend on the conditional law of the state process ${\mathcal{L}}(X^{\alpha}_{t}\,|\,{\mathscr{G}})$. We again derive a mean field game such that minimizers of the control problem give rise to Nash equilibria of this game.
We fix a probabilistic basis $\gamma$ and for given $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, $\alpha\in{\mathscr{A}}$, the state process $X^{\alpha}$ satisfies an equation of McKean-Vlasov type,
$\begin{split}X^{\alpha}_{t}=X_{0}+\int_{0}^{t}&b(s,X^{\alpha}_{\cdot\land s},\alpha_{s},\mu^{\alpha}_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha}_{\cdot\land s})\,\mathrm{d}{\mathscr{W}}_{s}+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha}_{\cdot\land s-},\zeta)\,\widetilde{\mathscr{N}}(\mathrm{d}s,\mathrm{d}\zeta),\end{split}$ (3.5)
where $\mu^{\alpha}_{t}={\mathcal{L}}(X^{\alpha}_{\cdot\land t}\,|\,{\mathscr{G}})$, $\boldsymbol{\sigma}=(\sigma^{0},\sigma)$, and $\boldsymbol{\lambda}=(\lambda^{0},\lambda)$. Under usual Lipschitz and linear growth assumptions, this equation has a unique strong solution for any $\alpha\in{\mathscr{A}}$, see Section 4.3. The mean field control problem is,
[…]
The main assumption that enables us to associate a mean field game to this control problem is the following (left) invertibility condition on the drift: If $B\in{\mathscr{B}}(\mathbb{R}^{n})$ denotes the image of the drift $b$, then there exists a function $\varphi:[0,T]\times{\mathscr{D}}_{n}\times B\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\to U$ such that $\varphi(t,x,b(t,x,a,\mu),\mu)=a$ holds for all $(t,x,a,\mu)\in[0,T]\times{\mathscr{D}}_{n}\times B\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})$. The precise assumptions are stated in Section 4.3, and a special case is presented in Section 5.3. Define running and terminal cost functions by,
$\begin{split}f(t,x,a,\nu)&=\int_{{\mathscr{D}}_{n}\times U}\delta_{\mu}F(t,\widetilde{x},\widetilde{a},\mu)(x)\,\nu(\mathrm{d}\widetilde{x},\mathrm{d}\widetilde{a})+F(t,x,a,\mu)\\ &\qquad+\int_{{\mathscr{D}}_{n}\times U}\nabla_{a}F(t,\widetilde{x},\widetilde{a},\mu)\cdot(\delta_{\mu}\varphi)(t,\widetilde{x},b(t,\widetilde{x},\widetilde{a},\mu),\mu)(x)\,\nu(\mathrm{d}\widetilde{x},\mathrm{d}\widetilde{a}),\\ g(x,\mu)&=\int_{{\mathscr{D}}_{n}}\delta_{\mu}G(\widetilde{x},\mu)(x)\,\mu(\mathrm{d}\widetilde{x})+G(x,\mu),\end{split}$
for $(t,x,a,\nu)\in[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ and as usual, $\mu=\Pi_{\sharp}\nu$. We emphasize that
$(\delta_{\mu}\varphi)(t,\widetilde{x},b(t,\widetilde{x},\widetilde{a},\mu),\mu)(x)=(\delta_{\mu}\varphi)(t,\widetilde{x},y,\mu)(x)\,\Bigl{|}_{y=b(t,\tilde{x},\tilde{a},\mu)}.$
For a given adapted measure flow $\mu=(\mu_{t})_{t\in[0,T]}$ on ${\mathscr{D}}_{n}$ and a control $\alpha\in{\mathscr{A}}$, we define the state process $X^{\alpha,\mu}$ of a representative player of the mean field game by,
$X^{\alpha,\mu}_{t}=X_{0}+\int_{0}^{t}b(s,X^{\alpha,\mu}_{\cdot\land s},\alpha_{s},\mu_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha,\mu}_{\cdot\land s})\,\mathrm{d}{\mathscr{W}}_{s}+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha,\mu}_{\cdot\land s-},\zeta)\,\widetilde{\mathscr{N}}(\mathrm{d}s,\mathrm{d}\zeta).$ (3.7)
The definition of a mean field game equilibrium reads analogous to Definition 3.3. We say that a pair $(\alpha^{*},\nu^{*})$ is a Nash equilibrium of the corresponding MFG of controls if the following hold:
  1. (i)
    The control $\alpha^{*}\in{\mathscr{A}}$ satisfies $J_{g}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J_{g}(X_{0},\alpha,\nu^{*})$ where, with an abuse of notation,
    $J_{g}(X_{0},\alpha,\nu)=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha,\mu}_{\cdot\land t},\alpha_{t},\nu_{t})\,\mathrm{d}t+g(X^{\alpha,\mu}_{\cdot\land t},\mu_{T})\right].$ (3.8)
    Here, $(f,g)$ are as in (3.4) and $X^{\alpha,\mu}$ solves (3.7) with $\mu=\Pi_{\sharp}\nu$.
  2. (ii)
[…]
In section 4.3 we prove the following result.
Theorem 3.7.
Suppose that Assumption 4.4 holds and that the MFC problem (3.6) is invariant under probability bases. Then, for $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, $\alpha^{*}\in{\mathscr{A}}$, and $\nu^{*}_{t}={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$,
$J(X_{0},\alpha^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J(X_{0},\alpha)\quad\Rightarrow\quad J_{g}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J_{g}(X_{0},\alpha,\nu^{*}),$
where $J$ is as in (3.6), and $J_{g}$ is as in (3.8).
Remark 3.8.
An inspection of our proof in Section 4.3 shows that the dependence on $(\alpha,\mu)$ could be through any of the coefficients $\sigma,\sigma^{0},\lambda,\lambda^{0}$ instead of the drift $b$ as in (3.5). Then, an analogous result to Theorem 3.7 holds provided the controlled coefficient satisfies an invertibility condition, and the other coefficients are independent of $(\alpha,\mu)$.
@@ 3.4 Law-Dependent Dynamics
The main idea is to compose the optimal control $\alpha^{*}$ with an arbitrary control $\alpha$ through a biased independent coin-flip. To achieve this, we extend the original probabilistic structure by adding an initial Bernoulli random variable, and then we use either one of the controls depending on its outcome. This provides linearity in the distribution variables and we obtain the result by essentially differentiating with respect to the success probability of the coin-flip. The assumed invariance principle ensures that $\alpha^{*}$ is still a minimizer of the extended problem.
@@ 4.1 Preliminaries
Then, for any $Z=Z(\omega,x)$ we have $Z=Z^{(0)}\otimes Z^{(1)}$ where $Z^{(x)}(\omega)=Z(\omega,x)$.
@@
This defines a family of probability bases $\gamma^{\delta}=(\hat{\Omega},\hat{\mathscr{F}},\mathbb{P}^{\delta},\hat{\mathbb{F}},\hat{\mathscr{W}},\hat{\mathscr{N}})$ indexed by $\delta\in(0,1)$. We note that $\hat{\mathscr{A}}:={\mathscr{A}}(\hat{\gamma})={\mathscr{A}}(\gamma^{\delta})$. For $\beta\in\hat{\mathscr{A}}$, the projected controls $\beta^{(0)},\beta^{(1)}$ are in ${\mathscr{A}}$ and by definition $\beta_{t}=\beta^{(0)}_{t}\otimes\beta^{(1)}_{t}$. Let $Y^{\beta}$ be the solution to $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$, and $X^{(0)}:=X^{\beta^{(0)}}$, $X^{(1)}:=X^{\beta^{(1)}}$ be the controlled state processes solving the equation SDE$(X_{0},\alpha)$ with controls $\beta^{(0)}$ and $\beta^{(1)}$, respectively. Further, we set,
@@
where the final equality follows from the construction of the stochastic integral. Hence, $I_{t}(\beta)=I_{t}(\alpha^{(0)})\otimes I_{t}(\alpha^{(1)})$ a.s. A similar computation of all other four integrals in $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$ implies that $X^{(0)}\otimes X^{(1)}$ is a solution of $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$. Then by strong uniqueness, we conclude that $Y^{\beta}=X^{(0)}\otimes X^{(1)}$.
Let $\xi:(\omega,x)\in\hat{\Omega}\to\xi(\omega,x)=x$ be the canonical projection. For $t\in[0,T]$ and a Borel set $S\subset{\mathscr{D}}_{n}\times U$, the followings hold a.s.,
$\nu^{\beta,\delta}_{t}(S)$ $=\mathbb{P}^{\delta}((Y^{\beta}_{\cdot\land t},\beta_{t})\in S\,|\,\hat{\mathscr{G}})$
@@
$\nu^{\beta,\delta}_{t}(S)(\omega,x)$ $=\mathbb{P}^{\delta}(\xi=0)\ \mathbb{P}((X^{(0)}_{\cdot\land t},\alpha^{(0)}_{t})\in S\,|\,{\mathscr{G}})+\mathbb{P}^{\delta}(\xi=1)\ \mathbb{P}((X^{(1)}_{\cdot\land t},\alpha^{(1)}_{t})\in S\,|\,{\mathscr{G}})$
$=\delta\,\nu^{(0)}_{t}(S)(\omega)+(1-\delta)\,\nu^{(1)}_{t}(S)(\omega).$
@@ 4.2 Completion of the Proof
$0\leq\lim_{\delta\downarrow 0}\frac{1}{\delta}\left(J_{\delta}(\hat{X}_{0},\beta)-J(X_{0},\alpha^{*})\right)=J_{g}(X_{0},\alpha,\nu^{*})-J_{g}(X_{0},\alpha^{*},\nu^{*}).$
In particular, $\alpha^{*}$ minimizes $J_{g}(X_{0},\alpha,\nu^{*})$ over all $\alpha\in{\mathscr{A}}$.
@@
$\hskip 50.0pt=\mathbb{E}\left[\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\land t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\,\Bigl{|}\,{\mathscr{G}}\,\right]\right]$
@@
We have shown that $\alpha^{*}$ minimizes $J_{g}(X_{0},\alpha,\nu^{*})$ and by its definition $\nu^{*}_{t}={\mathcal{L}}((X^{*}_{\wedge t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$. Hence, $(\alpha^{*},\nu^{*})$ is a Nash equilibrium of the MFG of controls.
@@
We use the invariance property only in (4.3). Therefore it suffices to require that the value function on $\gamma$ and its extension to $\hat{\gamma}$ coincide. In fact, if the probability basis $\gamma$ is rich enough in the sense defined in [23, Section 2.1.3] and $X_{0}$ is deterministic, one can directly prove that the extension to $\hat{\gamma}$ does not change the value.